OpenStudy (anonymous):

cannot figure out how to express the answer as multiples of pie 105 degree's being converted to radians

zepdrix (zepdrix):

\(\large\rm \pi\) radians on the unit circle is equivalent to \(\large\rm 180^o\).\[\large\rm \pi=180^o\]Dividing by 180 degrees gives us:\[\large\rm \frac{\pi}{180^o}=1\]So we can take our angle, and multiply it by this special 1,\[\large\rm 105^o\cdot\color{orangered}{1}=105^o\cdot\color{orangered}{\frac{\pi}{180^o}}\]Cancel out the units, degrees, from top and bottom,\[\large\rm =\frac{105\pi}{180}\]And simplify! :)